OpenStudy (anonymous):

simplify 1-e^x/1+e^x

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

simplify 1-e^x/1+e^x ( 1-e^x)/(1+e^x), add and subtract 1 =\[\frac{ -e ^{x}+1 }{ e ^{x}+1 }+\frac{ e ^{x}+1 }{ e ^{x}+1 }-1\] =\[\frac{ 2 }{ e ^{x}+1 }-1 \]